OpenStudy (anonymous):

The price of products may increase due to inflation and decrease due to depreciation. Derek is studying the change in the price of two products, A and B, over time. The price f(x), in dollars, of product A after x years is represented by the function below: f(x) = 0.69(1.03)x t (number of years) 1 2 3 4 f(t) (price in dollars) 10,100 10,201 10,303.01 10,406.04 Which product recorded a greater percentage change in price over the previous year?

OpenStudy (anonymous):

pleeeaseee somebodyy helpppp

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Product A has a fixed changing, that is 0.7107 while the changing of product B increases every year.

OpenStudy (anonymous):

How to get that? for A, just calculate f(x) when x =1,2,3,4, then take the difference between each year, you can see that it doesn't change. (the fixed equation gives us that conclusion also)

OpenStudy (anonymous):

for B, take the price of year 2 - the price of year 1, you get 10,201 - 10,100=101 then take the price of year 3- the price of year 2, you get 10,303.01-10201=102.01 (a little bit up, right?) do the same with year 4, 3 10,406.04-10,303.01=103.03 (a little bit up again, right?)

OpenStudy (anonymous):

so the price of product B increases year after year that allows me make the conclusion as above. :)

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (anonymous):

:)